I need help with all three parts. Thank you!

The following data represent the amount of time​ (in minutes) a random sample of eight students took to complete the online potion of an exam in a particular statistics course. Compute the​ mean, median, and mode time.

62.6
78.7
89.5
113.1
128.4
96.8
94.7
122.6

Compute the mean exam time. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.

A.

The mean exam time is

__?__ .

​(Round to two decimal places as​ needed.)

B.

The mean does not exist.

Compute the median exam time. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.

A.

The median exam time is __?__

​(Round to two decimal places as​ needed.)

B.

The median does not exist.

Compute the mode exam time. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.

A.

The mode is __?__.

​(Round to two decimal places as needed. Use a comma to separate answers as​ needed.)

B.

The mode does not exist.
Answer #1

data in ascending order : 62.6, 78.7, 89.5, 94.7, 96.8, 113.1, 122.6, 128.4

a) mean = (62.6 + 78.7 + 89.5 + 94.7 + 96.8 + 113.1 + 122.6 + 128.4) / 8 = 98.3

Option-A) The mean exam time is 98.3

b) median = middle most value = (94.7 + 96.8) / 2 = 95.75

Option-A) The median exam time is 95.75

c) Option-B) The mode does not exist